The average osmotic pressure of blood at 25°C is 7.7 atm. How many grams of glucose (C6H12O6) is needed to prepare a 110L solution that is isotonic with blood at 25°C?


1
Expert's answer
2021-01-15T07:01:29-0500

If the solution is isotonic, it means that it has the same osmotic pressure.

π=CRT

C=π/RT

T=25+273=298(K)

C=7.7/(0.082×298)=0.315(mol/L)

n=C×V=0.315×110=34.65(mol)

m=n×M

M=6×12+12×1+6×16=180(g/mol)

m=34.65×180=6237(g)